# IndiSeek learns information-guided disentangled representations

### v3

@@ A.4 MultiBench datasets
Table 4: MOSI results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@
Table 5: MOSEI results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@
Table 6: UR-FUNNY results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@
Table 7: MUStARD results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@
Table 8: MIMIC results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@
Table 9: Comparison of training time on multimodal datasets (max over $\lambda$, averaged over 10 seeds). All values are in seconds ($\times 10^{2}$).
@@ A.4.3 Ablation studies and comparison with CoMM
Table 10: MOSI ablation study results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
@@ A.4.4 Multi-task performance on MIMIC dataset
Table 11: Multi-task performance on MIMIC dataset (max over $\lambda$, averaged over 10 seeds). All values are percentages.

### v4

@@ A.4 MultiBench datasets
Table 4: MOSI results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@
Table 5: MOSEI results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@
Table 6: UR-FUNNY results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@
Table 7: MUStARD results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@
Table 8: MIMIC results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@
Table 9: Comparison of training time on multimodal datasets (averaged over 10 seeds, standard errors in parentheses, max average over $\lambda$). All values are in seconds ($\times 10^{2}$).
@@ A.4.3 Ablation studies and comparison with CoMM
Table 10: MOSI ablation study results with varying $\lambda$ (averaged over 10 seeds, standard errors in parentheses). All values are percentages.
@@ A.4.4 Multi-task performance on MIMIC dataset
Table 11: Multi-task performance on MIMIC dataset (averaged over 10 seeds, standard errors in parentheses, max average over $\lambda$). All values are percentages.